A nearsighted person wears contact lenses of refractive power -6.0 D. What is her far point?

17 cm

6.0 cm

-17 cm

-4.0 cm

-6.0 cm
Answer #1

focal length of the lens, f = -(1/P) = -(1/(-6)) m = (100/6) cm = 16.67 cm =17 cm (approx)

The far point of the person is 17.0 cm

Option (i) is the correct choice.